# Large Language Models, scientific knowledge and factuality: A systematic analysis in antibiotic discovery

### v2

@@ 4.1 Chemical compound definition generation
GPT-4: Top of the list and epistemically-aware (admits to lack of knowledge). 14/20 factual answers. For prompts without context, the model gave 70% (14/20) of the answers that were specific and with factuality. In two prompts (out of 20), the characterisation of a given chemical compound was non-factual due to the incorrect name of the fungus that may produce the chemical. Importantly, the model did not guess the definition if it did not recognize the name. For two chemical compounds, the model generated a request for more context or referred to the knowledge cutoff date of 2021-09. After adding context, GPT-4 generated text about the one chemical, but is factual for one out of two prompts. In the case of the second compound, the model still referred to the knowledge cutoff date of 2021-09. In general, adding context did not improve neither the factuality nor specificity metrics, which was still at 70%. The quoted names of the fungus that produces the chemical were incorrect. In general, the model generated fluent, semantically coherent, highly specific text (see Fumitremorgin C in Table S.2). We observed less hallucinations compared to ChatGPT.
@@
Llama 2: Exhibits a proclivity for generating inaccurate comparisons, particularly evident in chemical contexts. Llama 2-70B model produced 20% (4/20) responses that were both specific and factual. Smaller versions of Llama2 demonstrated inferior performance. Adding context failed to enhance either factuality or specificity metrics; instead, such augmentation resulted in a noteworthy decline, approaching zero, across all tested Llama 2 models. The Llama2 models displayed a noticeable tendency to generate comparative phrases, particularly evident when tasked with chemical comparisons, frequently resorting to statements like ’100 times more toxic than’ or ’100 times more potent than’. Specifically, the Llama2-7B produced such comparative statements in 70% of cases (14/20), while the Llama2-13B and Llama2-70B both did so in 40% of cases (8/20). These comparisons were consistently inaccurate, with a significant proportion mistakenly referencing morphine (10/30 such responses). In summary, the Llama 2 models failed to generate text characterized by fluency, semantic coherence, and high specificity.
@@
Table 3: Performance in Task1: chemical compound definition generation. Each step from the framework (see section 2.4) evaluated separately. X/20 is the number of correct responses vs total number of prompts. The bootstrapping result reported as % of total with 95% confidence interval in brackets.
@@
simple prompts P1-P3 context-based prompts P4-P7
@@
GPT-2 2/20 (0-30) 0/20 (0-0) 0/20 (0-0) 9/20 (10-70) 0/20 (0-0) 0-20 (0-0)
GPT-3 20/20 (100-100) 6/20 (0-60) 6/20 (0-60) 20/20 (100-100) 7/20 (10-60) 7/20 (10-60)
GPT-neo 8/20 (10-70) 0/20 (0-0) 0-20 (0-0) 11/20 (20-90) 0/20 (0-0) 0/20 (0-0)
Bloom 4/20 (0-40) 0/20 (0-0) 0/20 (0-0) 6/20 (10-60) 0/20 (0-0) 0/20 (0-0)
BioGPT 14/20 (40-100) 1/20 (0-20) 0/20 (0-0) 19/20 (80-100) 7/20 (10-60) 2/20 (0-30)
BioGPT-large 16/20 (50-100) 6/20 (0-60) 5/20 (0-50) 20/20 (100-100) 7/20 (10-60) 6/20 (0-60)
Galactica* 2/10 (0-50) 0/10 (0-0) 0/10 (0-0) 9/20 (70-100) 1/20 (0-30) 0/20 (0-0)
ChatGPT 20/20 (100-100) 12/20 (30-90) 12/20 (30-90) 20/20 (100-100) 11/20 (30-80) 10/20 (20-80)
GPT-4 20/20 (100-100) 14/20 (40-100) 14/20 (40-100) 20/20 (100-100) 14/20 (40-90) 14/20 (40-100)
Llama 2-7B 16/20 (50-100) 2/20 (0-30) 2/20 (0-30) 4/20 (0-40) 0/20 (0-0) 0/20 (0-0)
Llama 2-13B 12/20 (30-90) 2/20 (0-30) 2/20 (0-30) 9/20 (20-80) 0/20 (0-0) 0/20 (0-0)
Llama 2-70B 13/20 (40-90) 4/20 (0-50) 4/20 (0-50) 12/20 (30-90) 1/20 (0-20) 1/20 (0-20)
    *Evaluation for Galactica included P3 for simple prompts and P6-P7 for context-based prompts.
4.2 Chemical-fungus relation determination via entity generation
The aim of the task was to check whether the model generates a factual relation between the chemical compound and the fungus, which is producer of the given compound. We investigated ten compounds (five with and five without antibiotic activity), the same as in Task 1 described in section 3.1. We followed the proposed evaluation framework (Figure 3): first investigated the fluency, prompt-alignment and semantic coherence of the generated output, then the relation with generated fungus name and the factuality and specificity of the whole generated text. As prompts, we provided simple clausal fragments (e.g. ‘{entity 1} is isolated from’) or with an additional noun (e.g. ‘{entity 1} is produced by fungus’) as prefix (see Table 2). All models, except for Galactica, were evaluated via the 11 prompts designs. We report aggregated results for all 110 prompts (Table 4) and an optimal prompt with the best performance (Table 5). The optimal prompt configuration varies across models. In total we manually evaluated 1440 outputs in this task.
GTP-2, GPT-neo and Bloom fail to establish chemical-fungus relation. The GPT-2 model did not generate any answer containing the name of a potential fungus from which a given chemical compound can be isolated, regardless of the designed prompt. In addition, the model hallucinated new bacteria names (e.g. Streptococcus aureus). The GPT-neo model generated only 3.6% (4/110) of correct chemical compound-fungus relations. One correct answer each for Ergosterol, Alternariol, Atromentin and Flavasperone for three different prompts. The optimal prompt produced five chemical-fungus names relations, where only two relations were correct. Out of four correct chemical compound-fungus relations, only two had factual knowledge for the entire generated text (1.8% (2/110)).
The Bloom model generated only 2.7% (3/110) of factual chemical compound-fungus relations. In general, only 11% (12/110) of answers contained a fungus name (both factual and no factual relations). The model generated answers with the name of the fungus only for the two prompts, that hint towards fungus context: ‘{entity 1} is produced by the fungus’ and ‘{entity 1} is isolated from the fungus’. Considering the optimal prompt, Bloom produced seven fungi, two of them were in factual relation with the chemical. Out of three factual chemical compound-fungus relations, only two had factual knowledge for the entire generated text (1.8% (2/110)).
@@ 4.2 Chemical-fungus relation determination via entity generation
GPT-3 produces at least one factual relation for 6/10 compounds. GPT-3 generated 51% (56/110) answers containing a fungus name, where 23.6% (26/110) were factual compound-fungus relations, but only in 13.6% (15/110) the whole relation description was factual. The optimal prompt resulted in five factual relations, each containing individual fungus, of which only four answers were factual for whole description. The model produced a factual relation for at least one prompt for 6/10 tested compounds. Alternariol turned out to be a chemical with the most factual relations (8 out of 10 tested prompts), which aligns with the results from section 3.1 for GPT-3 (Table S.4).
@@
BioGPT biased towards Aspergillus, limited proportion of factual relations. BioGPT generated factual relations with fungi names for only 13.6% (15/110) of the cases, of which only ten answers had factual knowledge. Only three compounds were factually associated with the corresponding fungus from which they were isolated (Alternariol - six answers with factual relation, Atromentin - one answer with factual relation, Ergosterol - six answers with factual relation). Out of 62 generated answers that contained a fungus, 45 were Aspergillus, of which only 11 answers were factual. The two optimal prompts for this model were at 30% of the factual chemical compound-fungus relations.
@@
BioGPT-large: at least one factual chemical-fungus relation for 9/10 compounds. BioGPT-large generated 40.0% (44/110) of the factual chemical compound-fungus relations, of which less than half (18/44) were factual. The model generated at least one factual answer for 9 out of 10 tested compounds. This implies that it is possible to achieve 90% of factual relations if picking the right prompt for each chemical, showing the dependency on prompt optimisation. The model gave the most factual answers for Ergosterol (8 out of 10 tested prompts). If an optimal prompt is considered, BioGPT-large obtained 60% factual relations. The lack of factual knowledge resulted mainly from the incorrect name of the inhibitor of a given compound. Contrary to previous models, we observed that BioGPT-large did not generate additional context, when unclear (e.g. the names of the discoverers or the year of discovery of the compound if not known) reducing the overall number of hallucinations.
Galactica biased towards a large genus Aspergillus and poor performance. Galactica generated factual chemical-fungi relations for 6 out of 10 compounds. However, for four prompts we naively accepted a genus name Aspergillus as the factual answer, despite not being a specific name (Table S.5). Only for Fumitremorgin C the answer contained the species name: Aspergillus fumigatus. Aspergillus is at the same time the most frequently generated fungus (7 times). For Aphidicolin and Myriocin the model generated no answer. The generated answers consisted only of the name of the fungus, so no factual knowledge in the entire generated text was assessed.
@@
ChatGPT: coherent answers, regardless the prompt. ChatGPT generated 44.5% (49/110) factual chemical compound-fungus relations. The model provided all 11 fungal-name responses for eight compounds. The exception was one prompt for Conidiogenone and seven prompts for Ergosterol. The rest of the answers given for Ergosterol (4) were 100% factual with fungus relation. For three compounds all answers were factual (100% for Fumitremorgin C, Alternariol, Verrucarin A (Table S.6)) and for four compounds all prompts led to incorrect relations. The model produced very coherent answers regardless of the prompt, but only 31.8% (35/110) of them had a factuality for the whole generated description. The lack of factual knowledge was mainly due to the incorrect statement of the year of discovery/isolation of a given compound, the name of the discoverer or the class of the compound. The optimal prompt turned out to be ‘{entity 1} originally isolated from’, with 6/10 factual relations, of which the answer for only one relation had a factuality of whole answer. Almost 70% of the generated descriptions contained hallucinations.
@@
GPT-4 gives either the factual chemical-fungus relations or no answer at all. GPT-4 generated 50.9% (56/110) factual chemical compound-fungi relations. The model produced very coherent answers regardless of the prompt design. For four compounds, all answers were factual (11/11, 100% for Fumitremorgin C, Alternariol, Verrucarin A, Myriocin). The model for all prompts for Ergosterol did not generate any relations with the fungus. The model factually recognized the compound by citing its characteristics, but without giving the names of the fungi that produce ergosterol. In most cases, the model creates a new first sentence without completing the sentence from the prompt. Rarely in the first sentence is the name of the fungus that produces the chemical compound. In the case of 13.6% (15/110) responses, the model asked for more context and/or indicated that the question is incomplete. For 4.5% (5/110) prompts the model referred to the knowledge cutoff date of 2021-09, thus not providing an answer. Three prompts were optimal with 6/10 factual relations. Out of 56 factual chemical-fungi relations, the model generated 48 answers that had factual knowledge for the entire generated text (43.6%, 48/110). The hallucination (lack of factual knowledge) was observed for  56% of outputs, mainly as an incorrect statement of the year of discovery/isolation of a given compound or the incorrect citation of the name of the synonym of the fungus.
@@
Llama 2: the factuality increases with the model size. As the model size increases, the Llama 2 model exhibits a heightened capacity to generate fungal names and factual chemical compound-fungus relations, constituting 28.2% (31/110) of correct relations for Llama 2-70B. Notably, the Llama 2 70B accurately generated complete definitions, inclusive of the fungus name, with a success rate of 12.7% (14/110). Irrespective of the model size, Llama 2 exhibited a proclivity for response repetition and looping, with frequencies of 9% for Llama 2-7B, 4% for Llama 2-13B, and 5.5% for Llama 2-70B. Additionally, the model consistently supplied names of other chemical compounds as examples of fungi (24.5% Llama 2-7B, 34.5% Llama 2-13B, 27.3% Llama 2-70B) and occasionally substituted the expected fungal name with the year of compound discovery (13.6% Llama 2-7B, 12.7% Llama 2-13B, 25.5% Llama 2-70B). The Llama 2-7B and Llama 2-13B failed to provide fungal-name responses for the Ergosterol compound. Concerning the chemical compound Verrucarin A, all Llama 2 models erroneously postulated the name of a fungus (Verrucaria) unrelated to the actual producer of this compound (Albifimbria verrucaria or Myrothecium verrucaria), with instances recorded at 36.4% (4/11) for Llama 2-7B, 45.5% (5/11) for Llama 2-13B, and 36.4% (4/11) for Llama 2-70B. Prompt P18, structured as ’{entity 1} is produced by fungi, such as,’ yielded fungal names from all model variants for 9 out of 10 compounds. However, only Llama 2-13B and Llama 2-70B demonstrated factual relations, in 22.2% (2/9) and 55.6% (5/9), respectively.
Correctly generated relations are homogeneous and prevalent.
Considering the models with outputs that contain the fungus name, in majority they recognize one specific factual relation chemical-fungus, although the chemical can be produced by several fungi (RQ1). For instance, in all factually generated relations for Fumitremorgin C the fungus was Aspergillus fumigatus or Aspergillus (39 out of 39 correct fungi name, from 66 generated fungi in total). Other chemicals: Aphidicolin $\rightarrow$ Cephalosporium aphidicola (22/25 out of 55), Verrucarin A $\rightarrow$ Myrothecium verrucaria or Myrothecium (22/22 out of 73), (+)-Chloromonilicin $\rightarrow$ Monilinia fructicola or Monilinia (11/11 out of 53), Alternariol $\rightarrow$ Alternaria alternata or Alternaria fungus (61/67 out of 76), Flavasperone $\rightarrow$ Aspergillus (15/19 out of 39), Atromentin $\rightarrow$ Aspergillus (11/11 out ouf 54), Myriocin $\rightarrow$ Isaria sinclarii (17/21 out of 57). The least recognised chemical Conidiogenone was associated with Penicillium 9 out of 10 correct outputs (out of 55 outputs with fungi associations). Factually generated relations for Ergosterol are with two fungi: Aspergillus (11/23) and Ganoderma lucidum (7/23). Interestingly, Aspergillus is output by BioGPT and Ganoderma lucidum by BioGPT-Large. Although trained on the same corpus, we hypothesise that the higher number of parameters in BioGPT-Large (347M vs 1.5B) led to a more specific relation (RQ2).
@@
Table 4: Performance in Task 2: chemical-fungus relation determination via entity generation. Each step from the framework (see section 2.4) evaluated separately. X/110 is the number of correct responses vs total number of prompts. The bootstrapping result reported as % of total with 95% confidence interval in brackets.
@@
simple prompts + context-based prompts P8-P19
@@
GPT-2 0/110 (0-0) 0/110 (0-0) 0/110 (0-0) 0/110 (0-0) 0/110 (0-0)
GPT-3 57/110 (20-80) 56/110 (20-80) 26/110 (0-50) 15/110 (0-40) 15/110 (0-40)
GPT-neo 29/110 (0-60) 28/110 (0-50) 4/110 (0-20) 2/110 (0-10) 0/110 (0-0)
Bloom 13/110 (0-30) 12/110 (0-30) 3/110 (0-10) 2/110 (0-10) 2/110 (0-10)
BioGPT 62/110 (30-80) 57/110 (20-80) 15/110 (0-40) 10/110 (0-30) 2/110 (0-10)
BioGPT-large 84/110 (50-100) 74/110 (40-90) 44/110 (10-70) 18/110 (0-40) 14/110 (0-40)
Galactica* 9/10 (70-100) 8/10 (50-100) 6/10 (30-90) 0/10 (0) 0/10 (0)
ChatGPT 103/110 (70-100) 101/110 (70-100) 49/110 (20-70) 35/110 (10-60) 35/110 (10-60)
GPT-4 87/110 (50-100) 77/110 (40-100) 56/110 (20-80) 48 (10-70) 48 (10-70)
Llama 2-7B 69/110 (30-90) 36/110 (10-60) 7/110 (0-30) 2/110 (0-10) 2/110 (0-10)
Llama 2-13B 70/110 (30-90) 49/110 (10-80) 14/110 (0-40) 8/110 (0-30) 4/110 (0-20)
Llama 2-70B 76/110 (40-90) 62/110 (30-80) 31/110 (0-60) 14/110 (0-30) 14/110 (0-40)
    *Evaluation for Galactica included P19 prompt only.
Table 5: Performance for optimal prompts in Task 2: chemical-fungus relation determination via entity generation. For each model an optimal prompt was selected. Evaluation according to the framework (see Figure 3). X/10 is the number of correct responses vs total number of chemicals.

### v3

@@ 4.1 Task 1: Chemical compound definition generation
GPT-4: Top performer from the list and epistemically-aware (admits to lack of knowledge). 80% of factual answers for the best prompt (Table 3). The model did not guess the definition if it did not recognise the name. For two chemical compounds, the model generated a request for more context or referred to the knowledge cutoff date of 2021-09. In general, adding context did not improve neither the factuality nor specificity metrics, which remained at 70% (for two prompts in total, P1-P2 vs P3-P4). The quoted names of the fungus that produces the chemical were incorrect. In general, the model generated fluent, semantically coherent, highly specific text (Table A.4). We observed less hallucinations compared to ChatGPT.
@@
Llama 2: Exhibits a proclivity for generating inaccurate comparisons, particularly evident in chemical contexts. Llama 2-70B model produced 20% (4/20) responses that were both specific and factual (20% of factual answers for each prompt, P1 and P2). Smaller versions of Llama2 demonstrated inferior performance. Adding context failed to enhance either factuality or specificity metrics; instead, such augmentation resulted in a noteworthy decline, approaching zero, across all tested Llama 2 models. The Llama2 models displayed a noticeable tendency to generate comparative phrases, particularly evident when tasked with chemical comparisons, frequently resorting to statements like ‘100 times more toxic than’. In summary, the Llama 2 models failed to generate text characterized by fluency, semantic coherence, and high specificity.
@@
Table 2: Performance in Task1: chemical compound definition generation. Each step from the framework (see Section 3.1) evaluated separately. X/20 represents the number of correct responses versus the total number of outputs for a given type of prompt (each prompt has 10 outputs). The best results are highlighted in bold.
@@
simple prompts P1-P2 context-based prompts P3-P4
@@
GPT-2 2/20 0/20 0/20 9/20 0/20 0/20
GPT-3 20/20 6/20 6/20 20/20 7/20 7/20
GPT-neo 8/20 0/20 0/20 11/20 0/20 0/20
Bloom 4/20 0/20 0/20 6/20 0/20 0/20
BioGPT 14/20 1/20 0/20 19/20 7/20 2/20
BioGPT-large 16/20 6/20 5/20 20/20 7/20 6/20
ChatGPT 20/20 12/20 12/20 20/20 11/20 10/20
GPT-4 20/20 14/20 14/20 20/20 14/20 14/20
Llama 2-7B 16/20 2/20 2/20 4/20 0/20 0/20
Llama 2-13B 12/20 2/20 2/20 9/20 0/20 0/20
Llama 2-70B 13/20 4/20 4/20 12/20 1/20 1/20
Table 3: Performance for optimal prompts in Task 1: chemical compound definition generation. For each model an optimal prompt (highest score in STEP 2) was selected. Evaluation according to the framework (Fig. 1). X/10 is the number of correct responses vs total number of chemicals.
Model Optimal Prompt STEP1 STEP2 STEP 3
GPT-3 {entity 1} is a substance … 10/10 4/10 4/10
BioGPT {entity 1} is a compound … 9/10 5/10 1/10
BioGPT-large {entity 1} is a compound … 10/10 4/10 4/10
ChatGPT {entity 1} is a substance … 10/10 7/10 6/10
GPT-4 {entity 1} … 10/10 8/10 8/10
Llama 2-7B {entity 1} … 9/10 2/10 2/10
Llama 2-13B {entity 1} … 7/10 1/10 1/10
{entity 1} is … 5/10 1/10 1/10
Llama 2-70B {entity 1} … 7/10 2/10 2/10
{entity 1} is … 6/10 2/10 2/10
4.2 Task 2: Chemical-fungus relation determination via entity generation
@@
GPT-3 produces at least one factual relation for 6/10 compounds. GPT-3 generated 51% (56/110) answers containing a fungus name, where 23.6% (26/110) were factual compound-fungus relations, but only in 13.6% (15/110) the whole relation description was factual (STEP 2B). Alternariol turned out to be a chemical with the most factual relations (8 out of 10 tested prompts), which aligns with the results from section 3.1 for GPT-3 (Table A.5).
@@
BioGPT biased towards Aspergillus, limited proportion of factual relations. BioGPT generated factual relations with fungi names for only 13.6% (15/110) of the cases, of which only ten answers had factual knowledge. Out of 62 generated answers that contained a fungus, 45 were Aspergillus, of which only 11 answers were factual.
@@
BioGPT-large: at least one factual chemical-fungus relation for 9/10 compounds. This implies that it is possible to achieve 90% of factual relations if picking the right prompt for each chemical, showing the dependency on prompt optimisation. The model gave the most factual answers for Ergosterol (8 out of 10 tested prompts). Contrary to previous models, we observed that BioGPT-large did not generate additional context, when unclear (e.g. the names of the discoverers or the year of discovery of the compound if not known) reducing the overall number of hallucinations.
@@
ChatGPT: coherent answers, regardless of the prompt, but only 31.8% (35/110) of them were fully factual for the whole generated description (STEP 2B). The lack of factuality was mainly due to the incorrect statement of the year of discovery/isolation of a given compound, the name of the discoverer or the class of the compound. The model provided all 11 fungal-name responses for eight compounds. The exception was one prompt for Conidiogenone and seven prompts for Ergosterol. The rest of the answers given for Ergosterol (4) were 100% factual with fungus relation. For three compounds all answers were factual (100% for Fumitremorgin C, Alternariol, Verrucarin A (Table A.6)) and for four compounds all prompts led to incorrect relations.
@@
GPT-4 gives either the factual chemical-fungus relations or no answer at all. For four compounds, all answers were factual (11/11, 100% for Fumitremorgin C, Alternariol, Verrucarin A, Myriocin). The model for all prompts for Ergosterol did not generate any relations with the fungus and factually recognized the compound by citing its characteristics. In the case of 13.6% (15/110) responses, the model asked for more context and/or indicated that the question is incomplete. The hallucination (lack of factual knowledge) was observed for  56% of outputs, mainly as an incorrect statement of the year of discovery/isolation of a given compound or the incorrect citation of the name of the synonym of the fungus.
@@
Llama 2: the factuality increases with the model size. As the model size increases, the Llama 2 model exhibits an improved capacity to generate fungal names and factual chemical compound-fungus relations, constituting 28.2% (31/110) of correct relations for Llama 2-70B. Notably, the Llama 2 70B accurately generated complete definitions, inclusive of the fungus name, with a success rate of 12.7% (14/110). Irrespective of the model size, Llama 2 exhibited a proclivity for response repetition and looping. Concerning the chemical compound Verrucarin A, all Llama 2 models erroneously postulated the name of a fungus (Verrucaria) unrelated to the actual producer of this compound (Albifimbria verrucaria or Myrothecium verrucaria), with instances recorded at 36.4% (4/11) for Llama 2-7B, 45.5% (5/11) for Llama 2-13B, and 36.4% (4/11) for Llama 2-70B.
@@
Table 4: Performance in Task 2: chemical-fungus relation determination via entity generation. Each step from the framework (see Section 3.1) evaluated separately. X/110 represents the number of correct responses versus the total number of responses for a given model, with 110 derived from 11 prompts each applied to 10 chemicals. The best results are highlighted in bold.
@@
simple prompts + context-based prompts P5-P15
@@
GPT-2 0/110 0/110 0/110 0/110 0/110
GPT-3 57/110 56/110 26/110 15/110 15/110
GPT-neo 29/110 28/110 4/110 2/110 0/110
Bloom 13/110 12/110 3/110 2/110 2/110
BioGPT 62/110 57/110 15/110 10/110 2/110
BioGPT-large 84/110 74/110 44/110 18/110 14/110
ChatGPT 103/110 101/110 49/110 35/110 35/110
GPT-4 87/110 77/110 56/110 48/110 48/110
Llama 2-7B 69/110 36/110 7/110 2/110 2/110
Llama 2-13B 70/110 49/110 14/110 8/110 4/110
Llama 2-70B 76/110 62/110 31/110 14/110 14/110
Table 5: Performance for optimal prompts in Task 2: chemical-fungus relation determination via entity generation. For each model an optimal prompt (highest score in STEP 2B) was selected. Evaluation according to the framework (Fig. 1). X/10 is the number of correct responses vs total number of chemicals.